OpenStudy (anonymous):

Find the sum of all natural numbers lying between 100 and 1000 which are the multiples of 5

Directrix (directrix):

A number is divisible by 5 if the its last digit is a 0 or 5. There should be a neat way to do this without counting every single one of them. Let's think.

OpenStudy (anonymous):

@Directrix is it 98450 ?

Directrix (directrix):

How did you calculate the 98450?

OpenStudy (anonymous):

@Directrix t1=100 tn= 195 common difference(d) = 5 no of terms =[(tn-t1)/d]+1

Directrix (directrix):

Wow! I just reread the problem and saw that we are no longer counting but finding the sum. I need to start over.

Directrix (directrix):

I'm not following your nth term of the series.

OpenStudy (anonymous):

i got 99550 is it right ??

Directrix (directrix):

@aftabali --> Post your work on this thread, please.

OpenStudy (anonymous):

@Directrix S=\[(n/2)(2t1+(n-1)d)\]

OpenStudy (anonymous):

i think i might be wrong but work on it as a1=100 and l=1000 n=? d=5 putting these values into the formula l=a1+(n-1)d i got n=181 and by substituting this all in Sn=n/2{2a1+(n-1)d i got this 99550

Directrix (directrix):

I don't follow. Would you show the formula with the numbers you used?

OpenStudy (anonymous):

l=a1+(n-1)d => 1000=100+(n-1)(d) => 1000=100+5n-5 => 5n=905 and finally n=181 and by using formula Sn=n/2{2a+(n-1)d} we get Sn=181/5{2(100)+(181-1)(5)} => 181/2{2(100)+180(5)}=> 181{100+450}=> 99550 might this work

Directrix (directrix):

@JOYAL --> What do you think? To determine n. A_n = a_1 + (n-1) d 1000 = 100 + (n-1)5 900 = 5(n-1) 180 = n-1 181 = n Sum = (n)/2 (a_1 + a_n) Sum = [(181)/2]* ( 100 + 1000) Sum = 99 550

OpenStudy (anonymous):

@ directrix i did the same one well nice explanation